Please help me with this problem:
The sum of the roots of the equation x squared -20x + 9 = 0 is______________________
The product of its roots is____________________________
Thank you very much. Carrie

Found 2 solutions by user_dude2008, JimboP1977:
Answer by user_dude2008(1862)

ax^2+bx+c

sum of roots is 20 (-b/a)
product of roots is 9 (c/a)

Answer by JimboP1977(311)

x^2-20x+9=0
I tend to use completing the square as I find this the easiest method:
(x-10)^2 - 100 = -9
so x = 91^(1/2) + 10
x = -9.539 + 10 = 0.461
OR
x = 9.539 + 10 = 19.539
SO the sum of the roots = 0.461 + 19.539 = 20
Product of the roots = 9
Does that make sense?
Do you need further clarification?